Skip to main content
A-LevelMathematicsProbability and statistics9709/52 · Oct/Nov 2022 · Q57 Marks

Eric has three coins. One of the coins is fair. The other two coins are each biased so that the probability of obtaining a head on any throw is 1/4, independently of all other throws. Eric throws all three coins at the same time. Events A and B are defined as follows. A: all three coins show the same result B: at least one of the biased coins shows a head

  1. (a)Show that P(B) = 7/16.[2]
  2. (b)Find P(A | B).[2]
  3. (c)The random variable X is the number of heads obtained when Eric throws the three coins. Draw up the probability distribution table for X.[3]

Mark scheme

Marking points from the official mark scheme (Oct/Nov 2022). The mark or mark type is shown in brackets.

  1. Method 1: Scenarios identified ignoring unbiased coin
  2. [M1] All 3 different calculations seen unsimplified. P(BH1 BT2) = 1/4 × 3/4 = 3/16 P(BT1 BH2) = 3/4 × 1/4 = 3/16 P(BH1 BH2) = 1/4 × 1/4 = 1/16
  3. [A1] Clear identification of all scenarios, linked probabilities and sum. AG 3/16 + 3/16 + 1/16 = 7/16
  4. Method 2: Scenarios identified with all 3 coins
  5. [M1] All 6 different calculations seen unsimplified. P(H BH1 BT2) = 1/2 × 1/4 × 3/4 = 3/32 P(T BH1 BT2) = 1/2 × 1/4 × 3/4 = 3/32 P(H BT1 BH2) = 1/2 × 3/4 × 1/4 = 3/32 P(T BT1 BH2) = 1/2 × 3/4 × 1/4 = 3/32 P(H BH1 BH2) = 1/2 × 1/4 × 1/4 = 1/32 P(T BH1 BH2) = 1/2 × 1/4 × 1/4 = 1/32
  6. [A1] Clear identification of all scenarios, linked probabilities and sum. AG (1+3+3+1+3+3) / 32 = 14/32 = 7/16
  7. Method 3: 1- P(BT1 BT2) ignoring unbiased coin
  8. [M1] Calculation seen unsimplified and 1 - probability seen. 1 - (3/4)²
  9. [A1] Clear identification of scenario used, linked probability and calculation. AG 1 - 9/16 = 7/16
  10. Method 4: 1- P(BT1 BT2) with all 3 coins
  11. [M1] Both calculations seen unsimplified and 1 - 2 probabilities seen. 1 - P(H BT, BT2) - P(T BT, BT2) = 1 - (1/2 * 3/4 * 3/4) - (1/2 * 3/4 * 3/4)
  12. [A1] Clear identification of all scenarios used, linked probabilities and calculation. AG 1 - 9/32 - 9/32 = 1 - 18/32 = 14/32 = 7/16
  13. [M1] Their identified P(HHH) or correct as numerator and their identified P(B) or correct as denominator. Either numerical expression acceptable.
  14. [A1] Accept 0.071428... rounded to at least 3SF.
  15. [B1] Table with correct X values and at least one probability. Condone any additional X values if probability stated as 0.
  16. [B1] P(1) or P(2) correct, need not be in table, accept unsimplified.
  17. [B1] 4 correct probabilities linked with correct outcomes, may not be in table. Decimals correct to at least 3 SF. SC B1 for 4 probabilities (0 < p < 1) sum to 1 ± 0.005 with P(1) and P(2) incorrect.

Write your answer and get it marked

On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.

Sign up free →

About this question

Topic
Probability and statistics, A-Level Mathematics (syllabus 9709). Structured question, 7 marks.
Source
A-Level Mathematics, Oct/Nov 2022 Paper 5 Variant 2, question 5.

Questions and mark schemes are from past examination papers and remain the copyright of the exam board. How Oracle Prep builds its content

Related Mathematics questions

© 2026 Oracle Prep